OpenStudy (anonymous):

find equation of the line in slope intercept form: perpendicular to y=2x-3: (1,-2)

OpenStudy (anonymous):

y+2=-1/2(x-1)

OpenStudy (aroub):

y=mx+b , m=-1/2 , (1,-2) -2=-1/2(1)+b -2=-1/2+b 1/2+2=b 1/2+4/2=b 5/2=b the equation is y=-1/2x+5/2

OpenStudy (anonymous):

2y+4=-x+1 2y=-x-3

OpenStudy (anonymous):

aroub why are u using slope intercept form????

OpenStudy (aroub):

bec. 2 things one: i love to , two ; the question wants in the slope intercept form

OpenStudy (anonymous):

it says so in question, also, aroub, in the 4th line of your answer, why did the 2 suddenly become positive instead of negative?

OpenStudy (aroub):

lol ur so right!! i have no idea!

OpenStudy (anonymous):

but u r wrong

OpenStudy (aroub):

sorryy!!! :(

OpenStudy (anonymous):

lol, its okay, just try again? lol. I need the correct way to do this.

OpenStudy (anonymous):

y=-1/2x-3/2

OpenStudy (aroub):

okay sure yup as nerdsters answer! :)

OpenStudy (anonymous):

but how? like how do u get the 3/2? you divide the -2 from the y side?

OpenStudy (anonymous):

its 2

OpenStudy (anonymous):

nerdster, could you do the entire thing step by step? in one post?